### v1

# Ground states of strongly-correlated materials on quantum computers using ab initio downfolding

Abstract
The accurate first-principles description of strongly-correlated materials is an important and challenging problem in condensed matter physics. Ab initio downfolding has emerged as a way of deriving accurate many-body Hamiltonians including strong correlations, representing a subspace of interest of a material, using density functional theory calculations as a starting point. However, the solution of these material-specific models can scale exponentially on classical computers, constituting a challenge. Here we propose that utilizing quantum computers for obtaining the properties of downfolded Hamiltonians yields an accurate description of the ground state properties of strongly-correlated systems, while circumventing the exponential scaling problem. We demonstrate this for diverse strongly-correlated materials by combining ab initio downfolding and variational quantum eigensolvers, correctly predicting the antiferromagnetic state of one-dimensional cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic ground state of monolayer $\text{WTe}_{2}$, and the charge-ordered state of correlated metal $\text{SrVO}_{3}$. By utilizing a classical tensor network implementation of variational quantum eigensolvers we are able to simulate large models with up to $54$ qubits and encompassing up to four bands in the correlated subspace, which is indicative of the complexity that our framework can address.
I Introduction
Materials with strong electronic correlations exhibit a wealth of interesting properties, including superconductivity [1, 2], charge-ordered and spin-ordered states [3, 4], Mott insulating [5, 6] and excitonic insulating behavior [7, 8]. This constitutes strongly-correlated materials of great interest within condensed matter physics, and achieving a deep understanding of their properties central towards novel technological breakthroughs. It is therefore important to develop computational methods which can predict the properties of strongly-correlated systems from first quantum mechanical principles. Density functional theory (DFT) has been the main workhorse of computational materials science for decades [9], however it generally fails to capture strong electronic correlations, at least in its most commonly employed semilocal formulation [10].
A promising computational method for describing strongly-correlated systems from first principles is ab initio downfolding [11, 12, 13, 14], which utilizes a DFT starting point in order to derive a material-specific many-body Hamiltonian within a strongly-correlated active space of interest, which can be subsequently solved using a range of techniques, including exact diagonalization [15, 16], quantum Monte Carlo [17], diagrammatic approximations [18], and beyond. The many-body Hamiltonian resulting from downfolding and describing the physics of the strongly-correlated region is typically of the extended Hubbard form [19]:
$H=\sum_{\sigma}\sum_{\mathbf{R}\mathbf{R}^{\prime}}\sum_{ij}t_{i\mathbf{R}j\mathbf{R}^{\prime}}a_{i\mathbf{R}}^{\sigma\dagger}a_{j\mathbf{R}^{\prime}}^{\sigma}+$
$\frac{1}{2}\sum_{\sigma\rho}\sum_{\mathbf{R}\mathbf{R}^{\prime}}\sum_{ij}U_{i\mathbf{R}j\mathbf{R}^{\prime}}a_{i\mathbf{R}}^{\sigma\dagger}a_{j\mathbf{R}^{\prime}}^{\rho\dagger}a_{j\mathbf{R}^{\prime}}^{\rho}a_{i\mathbf{R}}^{\sigma},$ (1)
where $\sigma,\rho$ are spin indices, $\mathbf{R},\mathbf{R}^{\prime}$ denote lattice vectors, and $i,j$ run over the electronic bands of the system. Downfolding methods have been successfully applied to the description of charge-ordered systems and charge density waves [20, 21], high-temperature superconductors [13, 22, 23, 24, 25], and beyond. However, the unfavorable scaling of classical methods for obtaining eigenstates of the Hamiltonian of Eq. (I) has thus far prevented downfolding methods from being used for large systems, which can result in finite-size effects [26, 27], and limits the number of bands which may realistically be included in these models.
Quantum computers are a promising technology for the simulation of many-body quantum systems [28, 29], with a theoretical linear scaling of the computational cost with the number of orbitals, in an idealized scenario, given the 1:1 correspondence between qubits and spin-orbitals in models such as that of Eq. (I). This constitutes the simulation of strongly-correlated systems and Hubbard-like models an ideal problem for solution on quantum hardware [30, 31], and the performance of popular variational quantum algorithms in terms of producing an accurate ground state has recently been the subject of a detailed benchmark [32]. The potential of emerging fault-tolerant quantum computers to obtain properties of extended Hubbard models as that of Eq. (I) has been discussed in Ref. [33], where the necessary resources to access experimentally-relevant quantities have been estimated.
Here we demonstrate that current and near-term quantum computers can produce quantitatively and qualitatively accurate results on the strongly-correlated ground states of diverse materials. We do so by combining ab initio downfolding, in order to prepare material-specific Hamiltonians of the form of Eq. (I), with classical simulations of variational quantum eigensolver (VQE) optimizations to obtain the ground state of these Hamiltonians, utilizing a recently proposed tensor network implementation [34]. We show that our VQE energy for the ground state is quantitatively accurate when compared to results obtained within density matrix renormalization group (DMRG), and that we obtain the correct behavior for the ground state wavefunction and correlation functions across different scenarios, and specifically the antiferromagnetic behavior of the quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$ [35], the excitonic insulating ground state of the two-dimensional material $\text{WTe}_{2}$ [36, 7], and the charge-ordered state of the correlated metal $\text{SrVO}_{3}$ [37, 38]. These results, combined with the demonstrations that quantum simulation of downfolded Hamiltonians can yield accurate excited state properties in strongly correlated molecular systems [39, 40, 41], demonstrate the strong potential of current and emerging quantum computers for simulating the properties of strongly-correlated materials when combined with ab initio downfolding.
II Methods
II.1 Ab initio downfolding
The aim of ab initio downfolding approaches is to generate a many-body Hamiltonian representing an active space of interest for a given material, typically within the low-energy region in the vicinity of the Fermi level [11, 12, 13, 19], most often starting from a first-principles description such as DFT. An exchange term may be included in downfolded Hamiltonians [19], however we will ignore it here given its small magnitude for our studied systems, consistent with previous works [20]. The Hamiltonian of Eq. (I) obtained through ab initio downfolding contains hopping and Coulomb terms between all neighbors on a lattice, and indeed some of the longer-range terms, particularly next-nearest neighbor coupling, can be important in certain cases [42, 43]. Here we include nearest-neighbor terms only for the hopping and Coulomb terms, as these dominate over longer-range terms in the Wannier representation, due to the exponential decay of Wannier functions [44].
In order to describe periodic solids as a lattice model, we work within the basis of maximally-localized Wannier functions [45], which we will denote as $\phi$. Here the Wannier functions are centered at the different lattice sites $\mathbf{R}$ appearing in Eq. (I), with $i,j$ indexing the $i^{\text{th}}$ and $j^{\text{th}}$ Wannier function respectively. Starting from a Kohn-Sham DFT calculation performed within Quantum Espresso [46], the Wannier representation of the studied systems within the active space of interest is obtained using the wannier90 code [47]. The Wannierization procedure yields the hopping terms appearing in Eq. (I) as
[…]
where $H_{KS}$ the Kohn-Sham Hamiltonian and $V$ the volume of the unit cell. For $\mathbf{R}=\mathbf{R}^{\prime}$ this term represents the chemical potential of site $\mathbf{R}$. Moreover, following the wannierization using wannier90, we utilize the wan2respack [48] and RESPACK [19] software packages in order to obtain the two-body Coulomb integrals appearing in Eq. (I) as
$U_{i\mathbf{R}j\mathbf{R}^{\prime}}(\omega)=$
[…]
The Coulomb integral of Eq. (II.1) is frequency-dependent, however, it is a common approximation to take the static limit $\omega=0$ in the Hamiltonian of Eq. (I). While this approximation has been shown to lead to an over-screening of the Coulomb interactions [50], we will utilize it here, reserving a more rigorous treatment of dynamical effects for a future study, as these have been shown to be necessary to account for in order to achieve truly predictive accuracy for ground- and excited-state observables within ab initio downfolding [49, 50]. The screened Coulomb interaction $W$ appearing in the integral of Eq. (II.1), and by extension the Hubbard term $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$, is computed within the constrained random phase approximation (cRPA) [51]. The cRPA excludes the screening of the states within the active space, as their Coulomb interactions are explicitly included in the Hamiltonian of Eq. (I).
II.2 Tensor network VQE simulation
Our aim is to obtain the ground state of the Hamiltonian of Eq. (I), for studied materials with an active space represented by different sets of $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$ and $t_{i\mathbf{R}j\mathbf{R}^{\prime}}$ matrices. Before presenting our approach for computing the ground state, it is worth discussing some aspects of the complexity of this problem. For a Hubbard model on an $N_{x}\times N_{y}$ square lattice, with $N_{b}$ electronic bands included in the active space of the studied system, $n_{q}=2N_{x}N_{y}N_{b}$ is the number of qubits needed to simulate its properties, where the factor $2$ accounts for spin. Denoting $N=N_{x}N_{y}$, the number of terms appearing in the Hamiltonian of Eq. (I) is classified as follows, considering only nearest-neighbor contributions:
  • Intra-band hopping $t_{i\mathbf{R}i\mathbf{R}^{\prime}}$: The number of hopping terms within the same band is $2\times N_{b}\times(N_{x}(N_{y}-1)+N_{y}(N_{x}-1))$. This accounts for nearest-neighbor pairs along both horizontal and vertical directions.
  • Inter-band hopping $t_{i\mathbf{R}j\mathbf{R}}$: These are the hopping terms between different bands at the same site, resulting in $2\times N_{b}\times(N_{b}-1)\times N$ terms.
  • Chemical potential $t_{i\mathbf{R}i\mathbf{R}}$: For the on-site energies, there are $2\times N_{b}\times N$ terms.
  • On-site interaction $U_{i\mathbf{R}i\mathbf{R}}$ and $U_{i\mathbf{R}j\mathbf{R}}$: There are $N_{b}\times N$ on-site interaction terms for each band, with an additional $N_{b}\times(N_{b}-1)\times N$ terms for inter-band interactions.
  • Off-site interaction $U_{i\mathbf{R}i\mathbf{R}^{\prime}}$ and $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$: These terms include both intra-band and inter-band interactions between different sites, considering only nearest neighbors. The total is $N_{b}^{2}\times(N_{x}(N_{y}-1)+N_{y}(N_{x}-1))$.
  • Total number of terms:
    $N_{b}\times\bigg{[}(N_{b}+2)[N_{x}\times(N_{y}-1)+$
    $N_{y}\times(N_{x}-1)]+3N_{b}N\bigg{]}$
We do not include inter-band hopping between different sites to ensure a manageable computational cost for the systems studied here. Boundary effects have been implicitly considered by restricting hopping and off-site interactions to existing nearest-neighbor pairs within the lattice.
We simulate the VQE classically by representing the wavefunction as a matrix product state (MPS) [52] within a recently proposed variational tensor network eigensolver (VTNE) approach [34]. Specifically, following Ref. [34], we start from an initial state |$\ket{\psi_{o}}$, which is set to be the ground state of the non-interacting ($U=0$) case, and we apply a variational ansatz to generate a parameterized quantum state. We design a parameterized quantum circuit (PQC) as
|$\ket{\psi_{\mathrm{PQC}}(\bm{\theta})}=U_{n}(\bm{\theta}_{n})...U_{1}(\bm{\theta}_{1})\ket{\psi_{o}},$ (4)
where the precise form of the operators $U_{n}$ is determined by the choice of variational ansatz used in our simulations. Each of these operators takes as arguments a set of parameters $\bm{\theta}_{n}$, which are initialized randomly (or in some cases with a classical heuristic). The PQC is applied to a wave function which is represented as an MPS |$\ket{\psi_{\chi}(\bm{\theta})}$ with bond dimension $\chi$, and we can therefore compute the energy function
$E_{\chi}(\bm{\theta})=\bra{\psi_{\chi}(\bm{\theta})}H\ket{\psi_{\chi}(\bm{\theta})},$ (5)
with the Hamiltonian $H$ represented as a matrix product operator (MPO) [52]. Within our optimization scheme, we vary the parameters $\bm{\theta}$ of the PQC to minimize the energy in Eq. (5). In order to reduce the chance of becoming stuck in a local minimum, we perform ten independent optimizations, each with different random starting parameters $\bm{\theta}_{n}$, and we report the minimum energy in each case. This type of data analysis for exploring an energy surface with many local minima was performed in a previous study [32]. Additionally, following this energy minimization, we use the ten resulting approximate representations of the ground state for an optimization which instead minimizes the infidelity with respect to the DMRG solution, which we consider as the ground truth:
$\text{IF}=1-|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}.$ (6)
We find that this additional step leads to greatly improved energies and fidelities.
Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}$ the number of qubits. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy expectation value with respect to the exact PQC. For all tensor operations in this work we have used the ITensor software package [53]. We provide details on our optimization procedure in Section S1 of the Supplementary Information (SI). In preparing our PQC, we utilize two ansätze in this work, a number-preserving (NP) ansatz [31], which was designed specifically within the context of solving single-band Hubbard models, and a more generic excitation-preserving (EP) [54] ansatz, which allows us to straightforwardly prepare our circuits in the case of multi-band Hubbard models. The performance of both ansätze for producing accurate ground states has been discussed in Ref. [32]. We apply the ansätze to the solution of the non-interacting ($U=0$) Hubbard representation of our systems, prior to which we apply $R_{z}(\theta)$ gates to each qubit, which improves the optimization procedure.
It is worth emphasizing that the quantum simulation of materials using representations as the Hamiltonian of Eq. (I) has been the subject of detailed benchmarks, and found to likely be feasible with near-term resources, without necessitating fully fault-tolerant quantum computers [55]. Nevertheless, as we aim to build increasing complexity into the reduced Hamiltonians obtained through downfolding or related methods, and approach the full many-body limit, it might become necessary to utilize fault-tolerant architectures [33]. However, our work here is mostly agnostic to this.
With regards to the fault tolerant era, we perform a rough estimate of the resources needed for the ansatz considered here, and for this purpose it is interesting to estimate the number of $T$-gates which will be necessary in order to implement our circuits in a fault-tolerant fashion, where the one-qubit and two-qubit operators are approximated with a required accuracy $\epsilon$. As described previously, we apply an $R_{z}$ gate to each qubit prior to the application of the unitary operators associated with the variational ansatz. The approximation of a one-qubit gate has been found to be possible with $1.15\log_{2}(1/\epsilon)$ $T$-gates [56]. From this we have $n_{q}\times 1.15\log_{2}(1/\epsilon)$ $T$-gates associated with this step. Additionally, both the NP and EP ansätze involve the application of two-qubit gates with two parameters each. For a total of $n_{\text{params}}$ parameters in the variational ansatz we have $n_{\text{params}}/2$ two-qubit gates, each of which requires $\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates to represent [57]. Therefore, we have a total of $n_{q}\times 1.15\log_{2}(1/\epsilon)n+\frac{n_{\text{params}}}{2}\times\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates required for the fault-tolerant simulation of our systems with accuracy $\epsilon$.
III Results
Figure 1: Structures of the systems studied in this work; $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), $\text{WTe}_{2}$ (panel b), and $\text{SrVO}_{3}$ (panel c). Ca atoms are given in gray, Cu in blue, O in red, Sr in green, V in purple, W in silver and Te in gold.
system lattice size $N_{b}$ DMRG energy (eV) VQE energy (eV) Fidelity
$\text{Ca}_{2}\text{CuO}_{3}$ $10\times 1$ $1$ $-1.030$ $-1.019$ $99.8\%$
$\text{WTe}_{2}$ $2\times 2$ $4$ $55.103$ $55.225$ $90.5\%$
$\text{SrVO}_{3}$ $3\times 3$ $3$ $-116.192$ $-115.995$ $76.8\%$
Table 1: Studied systems, simulated lattice sizes, number of bands in the subspace, DMRG energies, and best values for the VQE energies and fidelities with respect to the DMRG solution. As outlined in the main text, we solve the Hubbard model representation of $\text{Ca}_{2}\text{CuO}_{3}$ at half-filling, of $\text{WTe}_{2}$ at full filling of the valence bands, and of $\text{SrVO}_{3}$ at half-filling of the lowest band.
Fig. 1 visualizes the structures of the systems we study here. These are chosen as demonstrate the capability of our approach to correctly predict the ground state properties of diverse strongly-correlated materials; $\text{Ca}_{2}\text{CuO}_{3}$ is a quasi-1D cuprate known to display antiferromagnetic behavior along the Cu-O chains [35]; $\text{WTe}_{2}$ has been proposed to host an excitonic ground state, i.e., one where correlated electron-hole pairs form spontaneously [36, 7]; and $\text{SrVO}_{3}$ is a correlated metal that exhibits substantial charge ordering [37, 38]. Table 1 summarizes the number of bands included in the active space of each system, the lattice size on which the downfolded Hamiltonian is solved, the DMRG energy, and the best value for the energy and fidelity obtained from the ten independent VQE optimizations we perform. For $\text{Ca}_{2}\text{CuO}_{3}$ we only perform a single optimization, as the simpler energy landscape of this system makes this sufficient for finding a low-energy, high-fidelity solution. More details on the optimization for each system are given below. In Section S1 of the SI we also give the VQE energy and fidelity values for all optimizations performed for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$. Moreover, we summarize details of all DFT and cRPA calculations employed for these systems in SI Section S2. We now discuss the properties of the ground state wavefunctions of the materials of Fig. 1 as obtained within our combined ab initio downfolding/VQE approach.
III.1 Antiferromagnetism in $\text{Ca}_{2}\text{CuO}_{3}$
Figure 2: Band structure of $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The spin-state occupation of the ground state of downfolded $\text{Ca}_{2}\text{CuO}_{3}$ (panel b) indicates antiferromagnetic behavior, with the dependence of the antiferromagnetic order parameter $m$ on the lattice size given in panel c.
Within the bulk structure of $\text{Ca}_{2}\text{CuO}_{3}$, Cu atoms form one-dimensional chains connected by O atoms, as seen in Fig. 1a, which results in well-known antiferromagnetic behavior. The band structure of this system is visualized in Fig. 2a, as obtained within DFT calculations at the PBE level [58]. We downfold the electronic structure of this system onto the active space consisting of the highest occupied bands, which has strong contributions from Cu $d$-orbitals. Wannier interpolation yields a band structure within this active space (red), which is in excellent agreement with the full DFT calculations. The derived Hubbard model parameters clearly demonstrate the 1D character of the system, with a dominant nearest-neighbor hopping term of $t=-0.491$ eV and nearest neighbor Coulomb repulsion $V=0.903$ eV along a single spatial direction. The on-site Coulomb repulsion was found to be $U=3.578$ eV, in good agreement with previous estimates [35].
We solve for the electronic ground state of $\text{Ca}_{2}\text{CuO}_{3}$, by performing VQE simulations of a one-dimensional, one-band Hubbard model using the above parameters, for a $10\times 1$ lattice at half-filling, i.e. we perform a $20$-qubit simulation of a Hamiltonian with $37$ terms ($20$ terms corresponding to the chemical potential of the single band are discarded since we can define $\mu=t_{i\mathbf{R}i\mathbf{R}}=0$). We use ten layers of the NP ansatz to represent our wavefunction, corresponding to $580$ variational parameters, which leads to a VQE energy within $11$ meV of the DMRG energy, and with a fidelity $\mathcal{F}=|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}=99.8\%$. Moreover, we see in Fig. 2b that the VQE ground state shows clear antiferromagnetic behavior, as is known to be the case for this material. We quantify the antiferromagnetic behavior by computing the staggered magnetization [59]
$m=\frac{1}{N_{x}N_{y}}\left|\sum_{x=1}^{N_{x}}\sum_{y=1}^{N_{y}}(-1)^{x+y}\langle\psi|\hat{S}_{z}(x,y)|\psi\rangle\right|,$ (7)
where |$\ket{\psi}$ the optimized VQE wavefunction. We visualize in Fig. 2c the value of $m$ for different lattices $N\times 1$, as a function of $N$, for $N=3-11$ (for $N=2$ no antiferromagnetic behavior is present). We see that $m$ assumes significant values, even if one were to extrapolate to $N\rightarrow\infty$, suggesting persistent antiferromagnetic behavior in this system in the thermodynamic limit.
III.2 Excitonic ground state in $\text{WTe}_{2}$
Figure 3: Band structure of $\text{WTe}_{2}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. Difference $-\delta n_{el}$ (panel b) in the total number of electrons within the bands of the strongly-correlated ground state within VQE and DMRG, from the number of electrons in the bands within the conventional band-insulating state. Positive values indicate hole formation, while negative values suggest excess electrons compared to the band insulating case.
Monolayer $\text{WTe}_{2}$ is a two-dimensional system, visualized in Fig. 1b, which has been proposed to be an excitonic insulator [36, 7], i.e., to host correlated electron-hole pairs in its ground state. The band structure of this system as obtained within semi-local DFT is visualized in Fig. 3a, in good agreement with previous reports [36]. Ref. [36] included a small amount of exact exchange in the DFT functional in order to induce a small gap around the Fermi level. Additionally, Ref. [60] demonstrated that a gap around the Fermi level opens when a static $GW$ approximation is used. In both cases, the subsequent solution of the so-called Bethe-Salpeter equation for excitonic states [61, 62] yields bound excitons of negative energy [36, 60], hence suggesting that the band insulator ground state predicted by DFT and/or $GW$ is unstable towards an excitonic one.
Here we downfold the electronic structure on the subspace of four bands around the Fermi level (red in Fig. 3a). We were not able to perfectly reproduce the Kohn-Sham band structure with our computed Wannier-interpolated bands, with small deviations persisting towards the band edges. However, since this region is far from the Fermi level, it does not affect the formation of bound excitons. The crossing of the second and third bands of the subspace near the Fermi level is immediately suggestive of the possibility of exciton formation. To verify this, we obtain the ground state of the downfolded four-band extended Hubbard model within this subspace, the parameters for which are given in SI Section S3. Given the insulating character of this material, we solve our Hamiltonian at full-filling of the lower two states for a $2\times 2$ lattice, corresponding to a $2\times 2\times 4\times 2=32$-qubit simulation of a Hamiltonian with $288$ terms. We limit the bond dimension to $\chi=512$, and we utilize twenty layers of the EP ansatz, which leads to $1,304$ variational optimization parameters. The obtained ground state energy from VQE is within $0.12$ eV of the DMRG value, while the ground state VQE wavefunction has a fidelity of $90.5\%$. The difference $-\delta n_{el}$ in the total number of electrons within the bands of the strongly-correlated ground state we obtain with VQE (and with DMRG), from the number of electrons in the bands within the conventional band-insulating state (i.e., full-filling of the two lower-lying bands within the subspace, while the two higher-lying bands are empty), is visualized in Fig. 3b. The first observation to make here is that the DMRG and the VQE predictions are in excellent agreement. Moreover, it becomes evident that in our reduced system of $16$ electrons ($4$ lattice sites, times $2$ spin directions, times $2$ fully occupied electronic bands when we initialize), a significant lack of electrons occurs in the two lower-lying bands, i.e., positively charged holes form in the ground state. At the same time, a very substantial electronic population resides in the bands above the Fermi level (i.e., in the “conduction” bands). This presence of electrons in the “conduction” bands and holes in the “valence” bands is a hallmark of excitonic insulating behavior. To further quantify the formation of exciton pairs, we compute the excitonic insulator order parameter [63]
$\Delta=\frac{U^{\prime}}{N_{x}N_{y}}\sum_{x,y,v,c,\sigma}\langle\psi|C^{\dagger}_{c,x,y,\sigma}C_{v,x,y,\sigma}|\psi\rangle$ (8)
where $v,c$ denote valence and conduction states respectively, and $U^{\prime}$ the inter-orbital, on-site Coulomb repulsion between electrons. Here we define $U^{\prime}$ as the average on-site, inter-orbital repulsive interaction between electrons, as these terms are given in SI Section S3. For a $2\times 2$ lattice we find $\Delta=0.693$, in reasonable agreement to the order parameter obtained using DMRG $\Delta_{DMRG}=0.876$ on the same lattice. The fact that VQE perfectly reproduces the DMRG band occupations, however leads to a finite error in the calculation of the order parameter is in agreement with previous observations on the Hubbard model, where it was found that even high-fidelity solutions can struggle to capture correlation functions [64], although our additional step of an overlap-based optimization substantially improves their accuracy [32]. Overall, for a small-sized lattice representation of $\text{WTe}_{2}$, our combined ab initio downfolding/VQE approach captures the previously reported excitonic ground state of this system, entirely from first principles.
III.3 Charge-ordered state in $\text{SrVO}_{3}$
Figure 4: Band structure of $\text{SrVO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The lattice site occupations in the first two bands of the $\text{SrVO}_{3}$ ground state (panel b) indicate a clear charge ordering.
The band structure of $\text{SrVO}_{3}$ is visualized in Fig. 4a. The Fermi level crosses the subspace of the three vanadium $d$-bands, and hence semi-local DFT suggests that this system is a regular metal. However, $\text{SrVO}_{3}$ is known to be a correlated metal and to exhibit charge order [37, 38]. In order to obtain these signatures of strong correlations with our VQE approach, we downfold the electronic structure onto the active space of the three bands crossing the Fermi level (visualized in red in Fig. 4a), and we obtain a three-band extended Hubbard model, the parameters for which are given in SI Section S3, and which are in close agreement to previous downfolding calculations for this system [48].
We perform VQE simulations for a $3\times 3$ lattice of this three-band Hubbard Hamiltonian with $423$ terms, corresponding to a simulation with $3\times 3\times 3\times 2=54$ qubits. We apply ten layers of the EP ansatz, corresponding to $1,168$ variational parameters, with a bond dimension of $\chi=512$. While we were unable to obtain a VQE ground state of this complex system with a fidelity higher than $76.8\%$, our solution has an energy which is only $0.2$ eV above that of the DMRG solution, and as we will see, reproduces key manifestations of strong electronic interactions. In order to quantify the charge ordering in the ground state wavefunction, we compute the charge disproportionation parameter
$\Phi=\frac{\left|\sum_{A}-\sum_{B}\right|}{N_{x}\cdot N_{y}},$ (9)
where $\sum_{A},\sum_{B}$ the total charge in two sublattices, which we define as $A$ and $B$ respectively. The total charge within a sublattice is obtained as $\sum_{A,B}=\sum_{b}\sum_{x}\sum_{y}\sum_{\sigma}n(x,y,b,\sigma)$, with $b$ a band index and a site belonging to sublattice $A$ if $(x+y)\mod 2=0$, and to sublattice $B$ otherwise. We find significant charge disproportionation with an order parameter $\Phi=0.30$, in reasonable agreement with DMRG simulations on the same lattice ($\Phi_{DMRG}=0.41$), and on a larger $4\times 4$ lattice ($\Phi_{DMRG}=0.40$). In Fig. 4b we visualize the charge distribution of the ground state wavefunction obtained with DMRG on a $4\times 4$ lattice, within the first two bands of the subspace, as this case most clearly illustrates the significant charge ordering which occurs. However, even for the $3\times 3$ lattices studied within VQE, we make qualitatively very similar observations.
IV Conclusions and outlook
In this work we have utilized a combined ab initio downfolding-VQE approach to describe the ground states of strongly-correlated materials. We apply our approach to the antiferromagnetic quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic insulator $\text{WTe}_{2}$, and the correlated metal $\text{SrVO}_{3}$, and we demonstrate that we correctly predict the key physics of these systems, while semi-quantitatively reproducing the ground state energy as obtained using DMRG. This demonstrates the potential of our approach to utilize current and emerging quantum computing technologies in order to accurately predict the ground-state properties of diverse strongly-correlated materials, entirely from first quantum mechanical principles.
Our work raises several questions and highlights the importance of pursuing various avenues of research. The variational approaches we employ here can struggle to quantitatively reproduce the true ground state energy in the case of systems downfolded on two-dimensional lattices, it will therefore be important to explore flexible ansätze as ADAPT-VQE [65] for such problems moving forward. This will be particularly important as we are moving towards the fault-tolerant era [33] and quantum computer architectures with more capabilities, making it possible to describe increasingly complex materials, using our and similar approaches. Moreover, while here we have employed a static approximation for the Coulomb interactions within the active space, it will be interesting to perform a rigorous treatment of the frequency-dependence, as has been described in recent works [49, 50]. Furthermore, here we have employed the DFT generalized gradient approximation as a starting point for downfolding; it will be interesting to explore the influence of the starting point moving forward, as it has been shown previously that optimal DFT starting points can result in improved descriptions of correlated states with higher levels of electronic structure calculations [66, 67]. Finally, while here we have accounted only for electronic screening on Coulomb interactions, lattice motions have recently been shown to be capable of substantially modifying the screening, particularly in polar materials [68, 69], which could lead to modified downfolded Hamiltonians and phase diagrams [70, 71]; we reserve a detailed discussion of these effects for a future study.
V Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0029167 and DDR-ERCAP0029710.