### v3

# Compressing Hamiltonians with ab initio downfolding for simulating strongly-correlated materials on quantum computers

Abstract
The accurate first-principles description of strongly-correlated materials is an important and challenging problem in condensed matter physics. Ab initio downfolding has emerged as a way of deriving compressed many-body Hamiltonians that maintain the essential physics of strongly-correlated materials. The solution of these material-specific models is still exponentially difficult to generate on classical computers, but quantum algorithms allow for a significant speed-up in obtaining the ground states of these compressed Hamiltonians. Here we demonstrate that utilizing quantum algorithms for obtaining the properties of downfolded Hamiltonians can indeed yield high-fidelity solutions. By combining ab initio downfolding and variational quantum eigensolvers, we correctly predict the antiferromagnetic state of one-dimensional cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic ground state of monolayer $\text{WTe}_{2}$, and the charge-ordered state of correlated metal $\text{SrVO}_{3}$. Numerical simulations utilizing a classical tensor network implementation of variational quantum eigensolvers allow us to simulate large models with up to $54$ qubits and encompassing up to four bands in the correlated subspace, which is indicative of the complexity that our framework can address. Through these methods we demonstrate the potential of classical pre-optimization and downfolding techniques for enabling efficient materials simulation using quantum algorithms.
I Introduction
Materials with strong electronic correlations exhibit a wealth of interesting properties, including superconductivity [1, 2], charge-ordered and spin-ordered states [3, 4], Mott insulating [5, 6] and excitonic insulating behavior [7, 8]. This constitutes strongly-correlated materials of great interest within condensed matter physics, and achieving a deep understanding of their properties central towards novel technological breakthroughs. It is therefore important to develop computational methods which can predict the properties of strongly-correlated systems from first quantum mechanical principles. Density functional theory (DFT) has been the main workhorse of computational materials science for decades [9], however it generally fails to capture strong electronic correlations, at least in its most commonly employed semilocal formulation [10].
The full many-body Hamiltonian describing a general material includes $M^{4}$ terms, where $M$ the number of orbitals. For realistic systems this can be prohibitively large. A promising computational method for generating a compressed representation of materials from first principles is ab initio downfolding [11, 12, 13, 14]. This technique utilizes a DFT starting point in order to derive a material-specific many-body Hamiltonian within a strongly-correlated active space of interest, which can be subsequently solved using a range of techniques, including exact diagonalization [15, 16, 17, 18], quantum Monte Carlo [19, 20, 21], diagrammatic approximations [22], and beyond. The many-body Hamiltonian resulting from downfolding and describing the physics of the strongly-correlated region is typically of the extended Hubbard form [23]:
$H=\sum_{\sigma}\sum_{\mathbf{R}\mathbf{R}^{\prime}}\sum_{ij}t_{i\mathbf{R}j\mathbf{R}^{\prime}}a_{i\mathbf{R}}^{\sigma\dagger}a_{j\mathbf{R}^{\prime}}^{\sigma}+$
$\frac{1}{2}\sum_{\sigma\rho}\sum_{\mathbf{R}\mathbf{R}^{\prime}}\sum_{ij}U_{i\mathbf{R}j\mathbf{R}^{\prime}}a_{i\mathbf{R}}^{\sigma\dagger}a_{j\mathbf{R}^{\prime}}^{\rho\dagger}a_{j\mathbf{R}^{\prime}}^{\rho}a_{i\mathbf{R}}^{\sigma},$ (1)
where $\sigma,\rho$ are spin indices, $\mathbf{R},\mathbf{R}^{\prime}$ denote lattice vectors, and $i,j$ run over the electronic bands of the system. Downfolding methods have been successfully applied to the description of charge-ordered systems and charge density waves [24, 25], high-temperature superconductors [13, 26, 27, 28, 29], and beyond. However, the unfavorable scaling of classical methods for obtaining eigenstates of the Hamiltonian of Eq. (I) has thus far prevented downfolding methods from being used for large systems, which can result in finite-size effects [30, 31, 32], and limits the number of bands which may realistically be included in these models.
Quantum computers are a promising technology for the simulation of many-body quantum systems [33, 34], especially with regards to the number of qubits needed to simulate Hamiltonians of the form of eq. (I), as many algorithms can utilize a 1:1 correspondence between the qubits and the spin-orbitals of the system. This constitutes the simulation of strongly-correlated materials and Hubbard-like models an ideal problem for solution on quantum hardware [35, 36], and the performance of popular variational quantum algorithms in terms of producing an accurate ground state has recently been the subject of a detailed benchmark [37]. The potential of emerging fault-tolerant quantum computers to obtain properties of extended Hubbard models as that of Eq. (I) has been discussed in Ref. [38], where the necessary resources to access experimentally-relevant quantities have been estimated.
The utility of quantum computers for materials simulation is currently limited by the fact that, as previously mentioned, the number of terms of the full Hamiltonian of a typical material scales as $M^{4}$. This can in turn lead to extremely high resource estimates in terms of gates and run times. Here we circumvent this limitation by demonstrating that current and near-term quantum algorithms can produce quantitatively and qualitatively accurate results for ground states of strongly-correlated materials, at a modest computational cost. We achieve this by utilizing compressed representations obtained via ab initio downfolding, and classical tensor networks simulations of the variational quantum eigensolver (VQE) [39]. We show that our VQE energy for the ground state is quantitatively accurate when compared to results obtained within density matrix renormalization group (DMRG), and that we obtain the correct behavior for the ground state wavefunction and correlation functions across different scenarios, and specifically the antiferromagnetic behavior of the quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$ [40], the excitonic insulating ground state of the two-dimensional material $\text{WTe}_{2}$ [41, 7], and the charge-ordered state of the correlated metal $\text{SrVO}_{3}$ [42, 43]. These results, combined with demonstrations that quantum simulation of downfolded Hamiltonians can yield accurate excited state properties in strongly correlated molecular systems [44, 45, 46], highlight the strong potential of current and emerging quantum computers for simulating the properties of strongly-correlated materials at a modest computational cost, when combined with ab initio downfolding.
Figure 1: Schematic illustration of ab initio downfolding. Starting from a low-level description of the full system of interest, an active space is identified, and a many-body lattice Hamiltonian is generated, representing the physics within the active space. A typical form for the downfolded Hamiltonian is the extended Hubbard form of eq. (I), where each lattice site can be associated with different hopping and Coulomb terms.
II Methods
II.1 Ab initio downfolding
The aim of ab initio downfolding approaches is to generate a many-body Hamiltonian on a lattice, representing an active space of interest for a given material, typically within the low-energy region in the vicinity of the Fermi level [11, 12, 13, 23]. This is schematically illustrated in Fig. 1. Downfolding often utilizes a first-principles description such as DFT as a starting point. An exchange term may be included in downfolded Hamiltonians [23], however we will ignore it here given its small magnitude for our studied systems, consistent with previous works [24]. In order to describe periodic solids as a lattice model, we work within the basis of maximally-localized Wannier functions [47], which we will denote as $\phi$. Here the Wannier functions are centered at the different lattice sites $\mathbf{R}$ appearing in Eq. (I), with $i,j$ indexing the $i^{\text{th}}$ and $j^{\text{th}}$ Wannier function respectively. Starting from a Kohn-Sham DFT calculation performed within Quantum Espresso [48], the Wannier representation of the studied systems within the active space of interest is obtained using the Wannier90 code [49]. The Wannierization procedure yields the hopping terms appearing in Eq. (I) as
[…]
where $H_{KS}$ the Kohn-Sham Hamiltonian and $V$ the volume of the unit cell. For $\mathbf{R}=\mathbf{R}^{\prime}$ this term represents the on-site potential of site $\mathbf{R}$. Moreover, following the wannierization using Wannier90, we utilize the wan2respack [50] and RESPACK [23] software packages in order to obtain the two-body Coulomb integrals appearing in Eq. (I) as
$U_{i\mathbf{R}j\mathbf{R}^{\prime}}(\omega)=$
[…]
The Coulomb integral of Eq. (II.1) is frequency-dependent, however, it is a common approximation to take the static limit $\omega=0$ in the Hamiltonian of Eq. (I). While this approximation has been shown to lead to an over-screening of the Coulomb interactions [55], we will utilize it here, reserving a more rigorous treatment of dynamical effects for a future study, as these have been shown to be necessary to account for in order to achieve truly predictive accuracy for ground- and excited-state observables within ab initio downfolding [51, 55]. The screened Coulomb interaction $W$ appearing in the integral of Eq. (II.1), and by extension the Hubbard term $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$, is computed within the constrained random phase approximation (cRPA) [56]. The cRPA excludes the screening of the states within the active space, as their Coulomb interactions are explicitly included in the Hamiltonian of Eq. (I).
II.2 Tensor network VQE simulation
We simulate the VQE classically by representing the wavefunction as a matrix product state (MPS) [57] within a recently proposed variational tensor network eigensolver (VTNE) approach [39]. Specifically, following Ref. [39], we start from an initial state |$\ket{\psi_{o}}$, which is set to be the ground state of the non-interacting ($U=0$) case, and we apply a variational ansatz to generate a parameterized quantum state. We design a parameterized quantum circuit (PQC) as
|$\ket{\psi_{\mathrm{PQC}}(\bm{\theta})}=U_{n}(\bm{\theta}_{n})...U_{1}(\bm{\theta}_{1})\ket{\psi_{o}},$ (4)
where the precise form of the operators $U_{n}$ is determined by the choice of variational ansatz used in our simulations. Each of these operators takes as arguments a set of parameters $\bm{\theta}_{n}$, which are initialized randomly (or in some cases with a classical heuristic). The PQC is applied to a wavefunction which is represented as an MPS |$\ket{\psi_{\chi}(\bm{\theta})}$ with bond dimension $\chi$, and we can therefore compute the energy function
$E_{\chi}(\bm{\theta})=\bra{\psi_{\chi}(\bm{\theta})}H\ket{\psi_{\chi}(\bm{\theta})},$ (5)
with the Hamiltonian $H$ represented as a matrix product operator (MPO) [57]. Within our optimization scheme, we vary the parameters $\bm{\theta}$ of the PQC to minimize the energy in Eq. (5). In order to reduce the chance of becoming stuck in a local minimum, we perform ten independent optimizations, each with different random starting parameters $\bm{\theta}_{n}$, and we report the minimum energy in each case. This type of data analysis for exploring an energy surface with many local minima was performed in a previous study [37]. Additionally, following this energy minimization, we use the ten resulting approximate representations of the ground state for an optimization which instead minimizes the infidelity with respect to the DMRG solution, which we consider as the ground truth:
$\text{IF}=1-|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}.$ (6)
We find that this additional step leads to greatly improved energies and fidelities.
Unless otherwise explicitly stated, we use the maximum bond dimension $\chi_{\rm max}=2^{n_{q}/2}$, where $n_{q}$ the number of qubits. Since this bond dimension is sufficient to exactly represent an arbitrary wavefunction on $n_{q}$ qubits, the MPS representation yields the energy expectation value with respect to the exact PQC. For all tensor operations in this work we have used the ITensor software package [58]. We provide details on our optimization procedure in Section A of the Appendix. In preparing our PQC, we utilize two ansätze in this work, a number-preserving (NP) ansatz [36], which was designed specifically within the context of solving single-band Hubbard models, and a more generic excitation-preserving (EP) [59] ansatz, which allows us to straightforwardly prepare our circuits in the case of multi-band Hubbard models. The performance of both ansätze for producing accurate ground states has been discussed in Ref. [37]. We apply the ansätze to the solution of the non-interacting ($U=0$) Hubbard representation of our systems, prior to which we apply $R_{z}(\theta)$ gates to each qubit, which improves the optimization procedure.
II.3 Hamiltonian compression and the measurement problem
Our aim is to obtain the ground state of the Hamiltonian of Eq. (I), for studied materials with an active space represented by different sets of $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$ and $t_{i\mathbf{R}j\mathbf{R}^{\prime}}$ matrices. Before presenting our results, it is worth discussing some aspects of the complexity of this problem. For a Hubbard model on an $N_{x}\times N_{y}$ square lattice, with $N_{b}$ electronic bands included in the active space of the studied system, $n_{q}=2N_{x}N_{y}N_{b}$ is the number of qubits needed to simulate its properties, where the factor $2$ accounts for spin. Denoting $N=N_{x}N_{y}$, the number of terms appearing in the Hamiltonian of Eq. (I) is classified as follows, considering only nearest-neighbor contributions:
  • Intra-band hopping $t_{i\mathbf{R}i\mathbf{R}^{\prime}}$: The number of hopping terms within the same band is $2\times N_{b}\times(N_{x}(N_{y}-1)+N_{y}(N_{x}-1))$. This accounts for nearest-neighbor pairs along both horizontal and vertical directions.
  • Inter-band hopping $t_{i\mathbf{R}j\mathbf{R}}$: These are the hopping terms between different bands at the same site, resulting in $2\times N_{b}\times(N_{b}-1)\times N$ terms.
  • On-site potential $t_{i\mathbf{R}i\mathbf{R}}$: For the on-site energies, there are $2\times N_{b}\times N$ terms.
  • On-site interaction $U_{i\mathbf{R}i\mathbf{R}}$ and $U_{i\mathbf{R}j\mathbf{R}}$: There are $N_{b}\times N$ on-site interaction terms for each band, with an additional $N_{b}\times(N_{b}-1)\times N$ terms for inter-band interactions.
  • Off-site interaction $U_{i\mathbf{R}i\mathbf{R}^{\prime}}$ and $U_{i\mathbf{R}j\mathbf{R}^{\prime}}$: These terms include both intra-band and inter-band interactions between different sites, considering only nearest neighbors. The total is $N_{b}^{2}\times(N_{x}(N_{y}-1)+N_{y}(N_{x}-1))$.
  • Total number of terms:
    $n_{\text{terms}}=N_{b}\times\bigg{[}(N_{b}+2)[N_{x}\times(N_{y}-1)+$
    $N_{y}\times(N_{x}-1)]+3N_{b}N\bigg{]}$
We do not include inter-band hopping between different sites to ensure a manageable computational cost for the systems studied here. Boundary effects have been implicitly considered by restricting hopping and off-site interactions to existing nearest-neighbor pairs within the lattice. It becomes clear that the compressed representation arising from ab initio downfolding, leads to a much improved scaling of $N_{b}^{2}N_{x}N_{y}$ to leading order, compared to the $(N_{b,f}N_{x}N_{y})^{4}$ scaling of the full many-body Hamiltonian. Note that here $N_{b,f}$ is the full number of bands of the system, and not only the ones within an active space, and the total number of terms appearing in the full-many body Hamiltonian is equal to $2(N_{x}N_{y}N_{b,f})^{2}+4N_{x}N_{y}N_{b,f})^{4}$. Therefore, the compression scales as $N_{b}^{2}/(N_{b,f}^{4}N_{x}^{3}N_{y}^{3})$. For the specific examples we study below, we precisely quantify the degree of Hamiltonian compression achieved through our downfolding procedure. The number of bands $N_{b,f}$ entering the full many-body Hamiltonian is set to be equal to the number of occupied bands, plus any empty states that enter the active space of the material within the downfolded representation. To also enable a more fair comparison, we give the number of terms of the full many-body Hamiltonian when it is restricted within the active space of the compressed representation.
The compressed representation of materials becomes particularly important when considering the problem of measuring expectation values of their Hamiltonian on quantum hardware. For a system described my $M$ qubits, depending on the specific measurement strategy, such as measuring qubit-wise commuting terms simultaneously, measuring a family of non-crossing (NC) terms simultaneously, or performing a basis rotation (BR) grouping, the lower bound for the number of measurements required in order to determine the energy for a single VQE energy evaluation scales as $M^{2}-M^{6}$ [60, 61]. Although a more detailed investigation is warranted, these lower bounds suggest that it is conceivable, when using methods such as NC and BR groupings with an idealized scaling close to $M^{2}$ for the number of measurements, to achieve accurate simulations of strongly-correlated materials on near-term quantum hardware. Even for less favorable scaling in real devices, the compressed representations of our systems (no more than $M=54$ qubits) could be particularly important in this direction.
Finally, it is worth highlighting that when performing a finite amount of measurements for different observables on real quantum hardware, the expectation values necessarily have a statistical uncertainty associated with them. Our classical tensor network simulation of VQE does not currently include shot noise as part of the simulation, but rather it represents the ultimate accuracy that could be expected from the simulation of the compressed material Hamiltonians.
II.4 Downfolded Hamiltonian simulation with near-term and fault-tolerant hardware
It is worth emphasizing that the quantum simulation of materials using representations as the Hamiltonian of Eq. (I) has been the subject of detailed benchmarks, and found to likely be feasible with near-term resources, without necessitating fully fault-tolerant quantum computers [61]. Indeed, as we outline in Section III for the different materials studied here, the number of qubits and two-qubit gates are well within the thresholds set in Ref. [61] as what might be feasible on near-term hardware. Moreover, as also detailed in Section III, by assuming that the fidelity of two-qubit gates is $99.9$[62], we find that for all studied systems here our circuit fidelities are in the range of $52-75$% thanks to our compression of the Hamiltonian, which places them well above the $10$% threshold that has been identified for obtaining useful results [63]. We emphasize that while this is only meant to provide a rough estimate, it clearly demonstrates the importance of the compression for reducing the measurement problem.
Nevertheless, as we aim to build increasing complexity into the reduced Hamiltonians obtained through downfolding or related methods, and approach the full many-body limit, it will become necessary to utilize fault-tolerant architectures [38]. However, our work here is mostly agnostic to this. We can obtain a rough estimate of the resources that would be needed for a fault-tolerant simulation of the downfolded Hamiltonians using the ansatz considered here, through computing different metrics. For example, one may compute the 1-norm of the Hamiltonian of eq. (I):
$||H||_{1}=\sum_{ij}\sum_{\mathbf{R}\mathbf{R}^{\prime}}|\tilde{t}_{i\mathbf{R}j\mathbf{R}^{\prime}}|+\frac{1}{2}\sum_{ij}\sum_{\mathbf{R}\mathbf{R}^{\prime}}|\tilde{U}_{i\mathbf{R}j\mathbf{R}^{\prime}}|,$ (7)
where $\tilde{t},\tilde{U}$ are dimensionless, obtained through dividing the hopping and Coulomb matrices by the value of the dominant intra-band nearest-neighbor hopping. Our definition for the 1-norm includes summations over all neighbors, despite the fact that within our VQE simulations we only include nearest neighbor terms. The localized nature of the Wannier functions results in terms beyond nearest neighbor being small, and our values for the 1-norm should be considered as an upper bound for the Hamiltonians we actually simulate. The 1-norm provides a bound for the Trotter error [64], and smaller values for this quantity are associated with more efficient quantum phase estimation [65].
For the purpose of fault-tolerant simulation it is also interesting to estimate the number of $T$-gates which will be necessary in order to implement our circuits, where the one-qubit and two-qubit operators are approximated with a required accuracy $\epsilon$. As described previously, we apply an $R_{z}$ gate to each qubit prior to the application of the unitary operators associated with the variational ansatz. The approximation of a one-qubit gate has been found to be possible with $1.15\log_{2}(1/\epsilon)$ $T$-gates [66]. From this we have $n_{q}\times 1.15\log_{2}(1/\epsilon)$ $T$-gates associated with this step. Additionally, both the NP and EP ansätze involve the application of two-qubit gates with two parameters each. For a total of $n_{\text{params}}$ parameters in the variational ansatz we have $n_{\text{params}}/2$ two-qubit gates, each of which requires $\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates to represent [67]. Therefore, we have a total of $n_{q}\times 1.15\log_{2}(1/\epsilon)n+\frac{n_{\text{params}}}{2}\times\mathcal{O}[16\log(1/\epsilon)+32]$ $T$-gates required for the fault-tolerant simulation of our systems with accuracy $\epsilon$.
III Results
Figure 2: Structures of the systems studied in this work; $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), $\text{WTe}_{2}$ (panel b), and $\text{SrVO}_{3}$ (panel c). Ca atoms are given in gray, Cu in blue, O in red, Sr in green, V in purple, W in silver and Te in gold.
system lattice size $N_{b}$ DMRG energy (eV) VQE energy (eV) Fidelity
$\text{Ca}_{2}\text{CuO}_{3}$ $10\times 1$ $1$ $6.005$ $6.028$ $99.3\%$
$\text{WTe}_{2}$ $2\times 2$ $4$ $115.029$ $115.097$ $96.2\%$
$\text{SrVO}_{3}$ $3\times 3$ $3$ $-105.383$ $-105.365$ $31.8\%$
Table 1: Studied systems, simulated lattice sizes, number of bands in the subspace, DMRG energies, and best VQE energies and fidelities with respect to the DMRG solution. As outlined in the main text, we solve the Hubbard model representation of $\text{Ca}_{2}\text{CuO}_{3}$ at half-filling, of $\text{WTe}_{2}$ at full filling of the valence bands, and of $\text{SrVO}_{3}$ at half-filling of the lowest band.
system $n_{q}$ $n_{2q,G}$ circuit fidelity $||H||_{1}$ $n_{\text{terms}}$
$\text{Ca}_{2}\text{CuO}_{3}$ $20$ $290$ $74.8$% $2.67\times 10^{2}$ $37$
$\text{WTe}_{2}$ $32$ $652$ $52.1$% $3.31\times 10^{2}$ $288$
$\text{SrVO}_{3}$ $54$ $484$ $55.8$% $2.315\times 10^{3}$ $423$
Table 2: Parameters relating to the near-term and fault-tolerant quantum simulation of the studied systems. Specifcally, we give the number of qubits ($n_{q}$) required to simulate the downfolded Hamiltonians, the number of two-qubit gates $n_{2q,G}$ in our circuits, and the respective circuit fidelity if one assumes a $99.9\%$ fidelity for the individual gates, as well as the 1-norm and the number of terms of the downfolded Hamiltonians.
Fig. 2 visualizes the structures of the systems we study here. These are chosen to demonstrate the capability of our approach to correctly predict the ground state properties of diverse strongly-correlated materials; $\text{Ca}_{2}\text{CuO}_{3}$ is a quasi-1D cuprate known to display antiferromagnetic behavior along the Cu-O chains [40]; $\text{WTe}_{2}$ has been proposed to host an excitonic ground state, i.e., one where correlated electron-hole pairs form spontaneously [41, 7]; and $\text{SrVO}_{3}$ is a correlated metal that exhibits substantial charge ordering [42, 43]. Table 1 summarizes the number of bands included in the active space of each system, the lattice size on which the downfolded Hamiltonian is solved, the DMRG energy, and the best value for the energy and fidelity obtained from the ten independent VQE optimizations we perform. For $\text{Ca}_{2}\text{CuO}_{3}$ we only perform a single optimization, as the simpler energy landscape of this system makes this sufficient for finding a low-energy, high-fidelity solution. More details on the optimization for each system are given below. In Section A of the Appendix we also give the VQE energy and fidelity values for all optimizations performed for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$. Moreover, we summarize details of all DFT and cRPA calculations employed for these systems in Section B of the Appendix. We now discuss the properties of the ground state wavefunctions of the materials of Fig. 2 as obtained within our combined ab initio downfolding/VQE approach.
III.1 Antiferromagnetism in $\text{Ca}_{2}\text{CuO}_{3}$
Figure 3: Band structure of $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The spin correlation function of the ground state of downfolded $\text{Ca}_{2}\text{CuO}_{3}$ (panel b) indicates clear antiferromagnetic behavior.
Within the bulk structure of $\text{Ca}_{2}\text{CuO}_{3}$, Cu atoms form one-dimensional chains connected by O atoms, as seen in Fig. 2a, which results in well-known antiferromagnetic behavior. The band structure of this system is visualized in Fig. 3a, as obtained within DFT calculations at the PBE level [68]. We downfold the electronic structure of this system onto the active space consisting of the highest occupied bands, which has strong contributions from Cu $d$-orbitals. Wannier interpolation yields a band structure within this active space (red), which is in excellent agreement with the full DFT calculations. The derived Hubbard model parameters clearly demonstrate the 1D character of the system, with a dominant nearest-neighbor hopping term of $t=-0.491$ eV and nearest neighbor Coulomb repulsion $V=0.903$ eV along a single spatial direction. The on-site Coulomb repulsion was found to be $U=3.578$ eV, in good agreement with previous estimates [40].
We solve for the electronic ground state of $\text{Ca}_{2}\text{CuO}_{3}$, by performing VQE simulations of a one-dimensional, one-band Hubbard model using the above parameters, for a $10\times 1$ lattice at half-filling, i.e. we perform a $20$-qubit simulation of a Hamiltonian with $37$ terms ($20$ terms corresponding to the on-site potential of the single band are discarded since we can define $t_{i\mathbf{R}i\mathbf{R}}=0$), compared to $2.83\times 10^{10}$ terms in the full many-body Hamiltonian, and $4.02\times 10^{4}$ terms in the many-body Hamiltonian, when restricted within the active space. We use ten layers of the NP ansatz to represent our wavefunction, corresponding to $580$ variational parameters, corresponding to $290$ two-qubit gates. If we assume a $99.9$% fidelity for two-qubit gates on near-term hardware, this suggests a $74.8$% circuit fidelity for a single VQE energy measurement. The 1-norm of the downfolded Hamiltonian in this system is found to be $||H||_{1}=2.67\times 10^{2}$, which is much lower than typical values of full many-body Hamiltonians [69], thus indicating the scalability of our approach to fault-tolerant architectures. These parameters, which relate to the near-term and fault-tolerant simulation of $\text{Ca}_{2}\text{CuO}_{3}$, and also of the other materials studied here, are summarized in Table 2.
Our tensor network simulation of VQE leads to a ground state energy within $23$ meV of the DMRG energy, and with a fidelity $\mathcal{F}=|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}=99.3\%$. Moreover, we compute the spin correlation function
$\langle C_{ij}\rangle=\langle S_{i}^{z}S_{j}^{z}\rangle-\langle S_{i}^{z}\rangle\langle S_{j}^{z}\rangle,$ (8)
and we plot $\langle C_{1j}\rangle$ in Fig. 3b. The alternating sign of the spin correlation function indicates clear antiferromagnetic behavior for $\text{Ca}_{2}\text{CuO}_{3}$, as expected, and we see that VQE is in near-perfect agreement to DMRG.
III.2 Excitonic ground state in $\text{WTe}_{2}$
Figure 4: Band structure of $\text{WTe}_{2}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. Difference $-\delta n_{el}$ (panel b) in the total number of electrons within the bands of the strongly-correlated ground state within VQE and DMRG, from the number of electrons in the bands within the conventional band-insulating state. Positive values indicate hole formation, while negative values suggest excess electrons compared to the band insulating case.
Monolayer $\text{WTe}_{2}$ is a two-dimensional system, visualized in Fig. 2b, which has been proposed to be an excitonic insulator [41, 7], i.e., to host correlated electron-hole pairs in its ground state. The band structure of this system as obtained within semi-local DFT is visualized in Fig. 4a, in good agreement with previous reports [41]. Ref. [41] included a small amount of exact exchange in the DFT functional in order to induce a small gap around the Fermi level. Additionally, Ref. [70] demonstrated that a gap around the Fermi level opens when a static $GW$ approximation is used. In both cases, the subsequent solution of the so-called Bethe-Salpeter equation for excitonic states [71, 72] yields bound excitons of negative energy [41, 70], hence suggesting that the band insulator ground state predicted by DFT and/or $GW$ is unstable towards an excitonic one.
Here we downfold the electronic structure on the subspace of four bands around the Fermi level (red in Fig. 4a). We were not able to perfectly reproduce the Kohn-Sham band structure with our computed Wannier-interpolated bands, with small deviations persisting towards the band edges. However, since this region is far from the Fermi level, it does not affect the formation of bound excitons. The crossing of the second and third bands of the subspace near the Fermi level is immediately suggestive of the possibility of exciton formation. To verify this, we obtain the ground state of the downfolded four-band extended Hubbard model within this subspace, the parameters for which are given in Section C of the Appendix. Given the insulating character of this material, we solve our Hamiltonian at full-filling of the lower two states for a $2\times 2$ lattice, corresponding to a $2\times 2\times 4\times 2=32$-qubit simulation of a Hamiltonian with $288$ terms, compared to $6.29\times 10^{8}$ terms in the full many-body Hamiltonian, and $2.63\times 10^{5}$ terms in the many-body Hamiltonian when restricted within the active space. We limit the bond dimension to $\chi=512$, and we utilize twenty layers of the EP ansatz, which leads to $1,304$ variational optimization parameters, corresponding to $652$ two-qubit gates and a $52.1$% circuit fidelity, if one were to assume a $99.9$% fidelity for the individual gates. The 1-norm of the downfolded Hamiltonian is $||H||_{1}=3.31\times 10^{2}$, which here too is significantly smaller than values of the order to $10^{4}$, which are obtained even for small chemical systems [69]. The obtained ground state energy from VQE is within $68$ meV of the DMRG value, while the ground state VQE wavefunction has a fidelity of $96.2\%$. The difference $-\delta n_{el}$ in the total number of electrons within the bands of the strongly-correlated ground state we obtain with VQE (and with DMRG), from the number of electrons in the bands within the conventional band-insulating state (i.e., full-filling of the two lower-lying bands within the subspace, while the two higher-lying bands are empty), is visualized in Fig. 4b. The first observation to make here is that the DMRG and the VQE predictions are in excellent agreement. Moreover, it becomes evident that in our reduced system of $16$ electrons ($4$ lattice sites, times $2$ spin directions, times $2$ fully occupied electronic bands when we initialize), a significant lack of electrons occurs in the two lower-lying bands, i.e., positively charged holes form in the ground state. At the same time, a very substantial electronic population resides in the bands above the Fermi level (i.e., in the “conduction” bands). This presence of electrons in the “conduction” bands and holes in the “valence” bands is a hallmark of excitonic insulating behavior. To further quantify the formation of exciton pairs, we compute the excitonic insulator order parameter [73]
$\Delta=\frac{U^{\prime}}{N_{x}N_{y}}\sum_{x,y,v,c,\sigma}\langle\psi|C^{\dagger}_{c,x,y,\sigma}C_{v,x,y,\sigma}|\psi\rangle$ (9)
where $v,c$ denote valence and conduction states respectively, and $U^{\prime}$ the inter-orbital, on-site Coulomb repulsion between electrons. Here we define $U^{\prime}$ as the average on-site, inter-orbital repulsive interaction between electrons, as these terms are given in Section C of the Appendix. For a $2\times 2$ lattice we find $\Delta=0.379$, which is underestimated compared to the order parameter obtained using DMRG $\Delta_{DMRG}=0.640$ on the same lattice. The fact that VQE perfectly reproduces the DMRG band occupations, however leads to a finite error in the calculation of the order parameter is in agreement with previous observations on the Hubbard model, where it was found that even high-fidelity solutions can struggle to capture correlation functions [74], although our additional step of an overlap-based optimization substantially improves their accuracy [37]. Overall, for a small-sized lattice representation of $\text{WTe}_{2}$, our combined ab initio downfolding/VQE approach captures the previously reported excitonic ground state of this system, entirely from first principles.
III.3 Charge-ordered state in $\text{SrVO}_{3}$
Figure 5: Band structure of $\text{SrVO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The lattice site occupations in the first two bands of the $\text{SrVO}_{3}$ ground state (panel b) indicate a clear charge ordering.
The band structure of $\text{SrVO}_{3}$ is visualized in Fig. 5a. The Fermi level crosses the subspace of the three vanadium $d$-bands, and hence semi-local DFT suggests that this system is a regular metal. However, $\text{SrVO}_{3}$ is known to be a correlated metal and to exhibit charge order [42, 43]. In order to obtain these signatures of strong correlations with our VQE approach, we downfold the electronic structure onto the active space of the three bands crossing the Fermi level (visualized in red in Fig. 5a), and we obtain a three-band extended Hubbard model, the parameters for which are given in Section C of the Appendix, and which are in close agreement to previous downfolding calculations for this system [50].
We perform VQE simulations for a $3\times 3$ lattice of this three-band Hubbard Hamiltonian with $423$ terms, compared to $7.34\times 10^{9}$ terms in the full many-body Hamiltonian, and $2.13\times 10^{6}$ terms in the many-body Hamiltonian, when restricted to the active space. In the compressed representation, our system is described using $3\times 3\times 3\times 2=54$ qubits. We apply ten layers of the EP ansatz, which requires to $1,168$ variational parameters, with a bond dimension of $\chi=512$. This corresponds to $584$ two-qubit gates, and assuming a fidelity of $99.9$% for each one, suggests a $55.8$% circuit fidelity on near-term hardware. The 1-norm of the downfolded Hamiltonian is $||H||_{1}=2.315\times 10^{3}$, which makes this system somewhat more challenging compared to the ones compared previously in terms of its potential simulation on fault-tolerant hardware. While we were unable to obtain a VQE ground state of this complex system with a fidelity higher than $31.8\%$, our solution has an energy which is only $18$ meV above that of the DMRG solution, and as we will see, qualitatively reproduces key manifestations of strong electronic interactions. In order to quantify the charge ordering in the ground state wavefunction, we compute the charge disproportionation parameter
$\Phi=\frac{\left|\sum_{A}-\sum_{B}\right|}{N_{x}\cdot N_{y}},$ (10)
where $\sum_{A},\sum_{B}$ the total charge in two sublattices, which we define as $A$ and $B$ respectively. The total charge within a sublattice is obtained as $\sum_{A,B}=\sum_{b}\sum_{x}\sum_{y}\sum_{\sigma}n(x,y,b,\sigma)$, with $b$ a band index and a site belonging to sublattice $A$ if $(x+y)\mod 2=0$, and to sublattice $B$ otherwise. We find significant charge disproportionation with an order parameter $\Phi=0.21$, which is however overestimated compared to DMRG simulations on the same lattice ($\Phi_{DMRG}=0.12$). In Fig. 5b we visualize the charge distribution of the ground state wavefunction obtained with DMRG on a $4\times 4$ lattice, within the first two bands of the subspace, as this case most clearly illustrates the significant charge ordering which occurs. However, even for the $3\times 3$ lattices studied within VQE, we make qualitatively very similar observations.
IV Conclusions and outlook
In this work we have utilized ab initio downfolding to generate compressed representations of strongly-correlated materials. We have shown that near-term quantum algorithms, such as the VQE, can produce the ground states of these compressed Hamiltonians and yields the expected behavior, at a modest computational cost and also in quantitative agreement with DMRG. We apply our approach to the antiferromagnetic quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic insulator $\text{WTe}_{2}$, and the correlated metal $\text{SrVO}_{3}$, and we correctly predict the key physics of these systems, while semi-quantitatively reproducing the ground state energy as obtained using DMRG. This highlights the potential of our approach to utilize current and emerging quantum computing technologies in order to accurately predict the ground-state properties of diverse strongly-correlated materials, entirely from first quantum mechanical principles. Additionally, our classical simulation of VQE may be used as a pre-optimization step to the simulation of strongly-correlated materials on quantum hardware, hence providing an excellent starting point and minimizing the quantum resource cost [39, 75].
Our work raises several questions and highlights the importance of pursuing various avenues of research. The variational approaches we employ here can struggle to quantitatively reproduce the true ground state energy in the case of systems downfolded on two-dimensional lattices, it will therefore be important to explore flexible ansätze as ADAPT-VQE [76, 77] for such problems moving forward. This will be particularly important as we are moving towards the fault-tolerant era [38] and quantum computer architectures with more capabilities, making it possible to describe increasingly complex materials, using our and similar approaches. Moreover, while here we have employed a static approximation for the Coulomb interactions within the active space, it will be interesting to perform a rigorous treatment of the frequency-dependence, as has been described in recent works [51, 55]. Furthermore, here we have employed the DFT generalized gradient approximation as a starting point for downfolding; it will be interesting to explore the influence of the starting point moving forward, as it has been shown previously that optimal DFT starting points can result in improved descriptions of correlated states with higher levels of electronic structure calculations [78, 79]. Finally, while here we have accounted only for electronic screening on Coulomb interactions, lattice motions have recently been shown to be capable of substantially modifying the screening, particularly in polar materials [80, 81], which could lead to modified downfolded Hamiltonians and phase diagrams [82, 83, 84, 85]; we reserve a detailed discussion of these effects for a future study.
V Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0029167 and DDR-ERCAP0029710.
Appendix A Details of VQE optimization
Following Ref. [39], we start by optimizing for the ground state of the non-interacting system ($U=0$). We use this state |$\ket{\psi_{o}}$ to generate the parameterized quantum circuit
|$\ket{\psi_{\mathrm{PQC}}(\bm{\theta})}=U_{n}(\bm{\theta}_{n})...U_{1}(\bm{\theta}_{1})\ket{\psi_{o}}$ (11)
as the starting point for the optimization of the interacting case, representing the real material of interest. The initial parameters $\bm{\theta}$ are obtained randomly from a Gaussian distribution $\mathcal{N}(0,10^{-5})$ with zero mean and variance $\sigma^{2}=10^{-5}$. We perform an energy minimization, until one of three conditions is met: the energy tolerance (defined by the absolute difference between the energy at the final step and the penultimate step) reaches $10^{-7}$, the energy gradient reaches $10^{-6}$, or the optimization reaches $500$ steps. The optimizations are performed using the L-BFGS method [86].
In order to obtain ground states with high fidelities, following the energy minimization described above, we use the resulting wave functions as a starting point for an overlap-based optimization, where the optimizer minimizes the loss function $f=\log_{10}(1-|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2})$ (the logarithm of the infidelity with respect to the DMRG ground state). We find that this hybrid optimization strategy yields the overall best results in terms of minimizing the energy and also producing a high-fidelity ground state. We perform ten independent optimizations following the above hybrid energy-/overlap-based minimization, and we take the state with the minimal energy among these ten as the ground state, which prevents the system from becoming stuck in local minima.
An important factor in the ground state optimization is the electron filling of the bands of the different materials studied here. $\text{Ca}_{2}\text{CuO}_{3}$ and $\text{SrVO}_{3}$ are metallic at the DFT level of theory, and we solve for their ground state at half-filling of the first electronic band included in the model. $\text{WTe}_{2}$ is predicted to be a conventional band insulator at the Kohn-Sham level of theory, we therefore solve for the ground state of this system with an initial state where the two lower-energy bands in the active space (see Fig. 3 in main manuscript) are at full-filling, and the two upper bands are empty. Since here we work in the basis of Wannier functions, we populate the Wannier states with the greatest contribution from the two lower-energy Kohn-Sham states, averaged across the Brillouin zone, as can be deduced from the Wannier rotation matrices obtained within Wannier90 [49].
In Tables 3-6 we give the VQE energies and fidelities obtained for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$, ranked from best to worst from right to left.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$115.097$ eV $117.327$ eV $118.903$ eV $122.972$ eV $129.077$ eV $130.034$ eV $131.099$ eV $133.006$ eV $134.170$ eV $134.216$ eV
DMRG Energy: $115.029$ eV
Table 3: VQE energies of $\text{WTe}_{2}$.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$96.2\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Table 4: VQE fidelities of $\text{WTe}_{2}$ with respect to DMRG reference.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$-105.365$ eV $-105.363$ eV $-105.298$ eV $-105.260$ eV $-105.167$ eV $-105.120$ eV $-101.833$ eV $-84.214$ eV $-84.094$ eV $-82.473$ eV
DMRG Energy: $-105.383$ eV
Table 5: VQE energies of $\text{SrVO}_{3}$.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$31.8\%$ $11.7\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Table 6: VQE fidelities of $\text{SrVO}_{3}$ with respect to DMRG reference.
Appendix B Details of DFT and constrained RPA calculations
We perform all DFT calculations within the Quantum Espresso software package [48], within the generalized gradient approximation (GGA) of Perdew, Burke and Ernzerhof (PBE) [68]. We utilize scalar-relativistic optimized norm-conserving Vanderbilt pseudopotentials (ONCV) [87] with standard accuracy, as these are given in Pseudo Dojo [88].
For our DFT calculations on $\text{Ca}_{2}\text{CuO}_{3}$ we employ a wave function cutoff of $80$ Ry, and a $6\times 6\times 6$ k-grid. We compute the dielectric function and the Coulomb integrals of the system within RESPACK [23], using a polarizability cutoff of $7$ Ry and $100$ bands, excluding a single band crossing the Fermi level within cRPA. For $\text{WTe}_{2}$ we use a wave function cutoff of $80$ Ry, and a $6\times 6\times 1$ k-grid. We compute the dielectric function within cRPA by excluding the four bands around the Fermi level, with a polarizability cutoff of $5$ Ry and $600$ bands. For $\text{SrVO}_{3}$ we use a wave function cutoff of $81$ Ry, and a $6\times 6\times 6$ k-grid. We compute the dielectric function within cRPA by excluding the three bands crossing the Fermi level, with a polarizability cutoff of $5$ Ry and $600$ bands, yielding Coulomb parameters in close agreement to those reported previously [23].
Appendix C Downfolded Hamiltonian parameters
Here we give the parameters resulting from downfolding the electronic structure of the different materials onto the extended Hubbard Hamiltonian of eq. 1 of the main manuscript. Here we give the nearest-neighbor terms along the crystallographic direction where the maximal coupling and interactions occur, for each system.
C.1 $\text{Ca}_{2}\text{CuO}_{3}$
The hopping and Coulomb terms of this system are dominant along the crystallographic direction which aligns with chains of Cu atoms. We therefore construct a one-dimensional, single-band Hubbard model with the following parameters resulting from the Wannierization and downfolding procedures: hopping integral of $t=-0.491$ eV, on-site Coulomb interaction of $U=3.578$ eV and off-site Coulomb repulsion of $V=0.903$ eV.
C.2 $\text{WTe}_{2}$
Within the subspace of four bands around the Fermi level of $\text{WTe}_{2}$, we find the hopping term (all values in eV)
$t_{i\mathbf{R}j\mathbf{R}^{\prime}}=\begin{pmatrix}-0.201&0.178&-0.398&-0.128\\ 0.108&-0.144&0.072&-0.071\\ 0.398&0.003&0.387&0.025\\ 0.019&0.071&0.057&0.124\\ \end{pmatrix}.$ (12)
with $i=j$ the intra-band terms, and $i\neq j$ the inter-band contributions, for $\mathbf{R},\mathbf{R}^{\prime}$ corresponding to nearest neighbors.
Similarly, for the on-site Coulomb interaction
$U_{ij}=\begin{pmatrix}1.107&0.822&0.922&0.765\\ 0.822&1.095&0.760&0.684\\ 0.922&0.760&1.096&0.853\\ 0.765&0.684&0.853&1.174\\ \end{pmatrix},$ (13)
and the nearest-neighbor off-site terms:
$V_{ij}=\begin{pmatrix}0.924&0.822&0.841&0.765\\ 0.754&0.917&0.715&0.672\\ 0.841&0.760&0.855&0.853\\ 0.721&0.672&0.762&0.860\\ \end{pmatrix}.$ (14)
C.3 $\text{SrVO}_{3}$
$\text{SrVO}_{3}$ has cubic symmetry, making the Hamiltonian parameters identical along the three crystallographic axes. We find within the subspace of the three electronic bands crossing the Fermi level that we have the following intra- and inter-band terms, where all values are given in eV. For the hopping term
$t_{i\mathbf{R}j\mathbf{R}^{\prime}}=\begin{pmatrix}-0.263&0&0\\ 0&-0.263&0\\ 0&0&-0.027\\ \end{pmatrix}.$ (15)
The on-site Coulomb interaction
$U_{ij}=\begin{pmatrix}3.527&2.349&2.349\\ 2.349&3.527&2.349\\ 2.349&2.349&3.527\\ \end{pmatrix},$ (16)
and the nearest-neighbor off-site terms:
$V_{ij}=\begin{pmatrix}0.649&0.635&0.555\\ 0.635&0.649&0.555\\ 0.555&0.555&0.492\\ \end{pmatrix},$ (17)
where here too all values are given in eV.